### v1

# Pathspace Kalman Filters with Dynamic Process Uncertainty for Analyzing Time-course Data

Abstract
Kalman Filter (KF) is an optimal linear state prediction algorithm, with applications in fields as diverse as engineering, economics, robotics, and space exploration. Here, we develop an extension of the KF, called a Pathspace Kalman Filter (PKF) which allows us to a) dynamically track the uncertainties associated with the underlying data and prior knowledge, and b) take as input an entire trajectory and an underlying mechanistic model, and using a Bayesian methodology quantify the different sources of uncertainty. An application of this algorithm is to automatically detect temporal windows where the internal mechanistic model deviates from the data in a time-dependent manner. First, we provide theorems characterizing the convergence of the PKF algorithm. Then, we numerically demonstrate that the PKF outperforms conventional KF methods on a synthetic dataset lowering the mean-squared-error by several orders of magnitude. Finally, we apply this method to biological time-course dataset involving over 1.8 million gene expression measurements.
1 Introduction
Estimating sources of uncertainty about a system from data is a quintessential problem with applications in many branches of science and engineering [9], [42]. The Kalman Filter (KF) is a highly successful and provably optimal solution to this problem under certain assumptions [1]. Crucially, KFs carefully account for different sources of noise, including data variance, model variance, and model confidence (called process uncertainty) in order to yield optimal estimates. Figure 1A illustrates the key components of a KF, and Figure 1B illustrates the dynamic state prediction of a standard KF. The KF formalism is quite flexible and allows diverse domain-specific internal models coupled with a general numerically tractable estimation procedure. This flexibility of KFs has allowed for its successful application in diverse fields such as weather prediction [16], target tracking [17], data smoothing and assimilation [6], flight [25], parameter estimation in electro-chemical responses [3], microscopy [18], image processing [46], economics [4], music retrieval [29], and more.
More generally, KFs can be seen as simple probabilistic graphical models [14]. In this light, the KF models a system as a normal distribution with mean and variance derived as a convex combination of a measurement normal distribution and the internal model normal distribution, (Figure 1A) allowing the relative contributions of noise from these sources to be quantified. Kalman Filters are traditionally used to iteratively predict the state of a system at a future time given a stream of incoming measurements (Figure 1B). In many time-course experiments where KFs are used as an analysis tool, the entire measurement trajectory may be collected prior to analysis [22], [26, 30], [11, 20].
However, most methods fail to make use of the entire data trajectory simultaneously when optimizing the KF and instead iteratively run the filter along the data forward in time. Any procedure that fails to capitalize on the availability of the entire time trajectory may result in less than optimal estimates if future data may somehow constrain earlier predictions. Moreover, in a context with dynamical data and a potentially time-varying generative process, as is common in biology, there should be no apriori assumptions about the confidence in the underlying model being equal at all times times. Rather, we propose estimating the process uncertainty at different times in order to dynamically quantify the reliability of the internal model. For example, if there is a change in the underlying data generation process which is not accounted for in the internal model, the process uncertainty should adapt to this discrepancy. These observations motivate the development of a KF which uses information from an entire trajectory iteratively (Figure (1C)) to improve its estimates, including dynamically estimating process uncertainty, as opposed to only iterating forward in time.
Contributions Summary: We make the following contributions to the existing standard Kalman Filter algorithm:
1. Pathspace Kalman Filter (PKF): In Section 3.1, we describe the Pathspace Kalman Filter, which analyzes the entire time-series at once and dynamically quantifies process uncertainty. We prove the PKF allows for non-monotonic process uncertainty, yet that the PKF algorithm converges.
2. Bayesian Computation: We describe a scalable Bayesian methodology for simulation based computation of model expectation and variance in Section 4. This allows the PKF to be run efficiently on large datasets such as transcriptomics data.
3. Synthetic Data Experiments: In Section 5 we compare a standard KF and our PKF models on synthetic data and show a decrease in mean-squared-error by several orders of magnitude.
4. Real Data Experiments: In Section 5.2 we apply the PKF to a large time-course gene expression dataset with over 1.8 million measurements to demonstrate the utility and scalability of our method in the context of biological data.
Figure 1: A. Illustration of the Kalman filtering framework. A mechanistic model produces a distribution of predictions which are combined with the underlying data and process uncertainty to produce a low variance state estimate (filter prediction). B. Illustration of a standard Kalman filter which makes predictions forward in time as new data is acquired. C. Illustration of a pathspace Kalman filter which ingests the entire temporal trajectory to produce a an entire trajectory as output.
2 Background
2.1 Kalman Filters
A 1-dimensional Kalman Filter (KF) is an optimal prediction-measurement-update algorithm in the sense that given a linear system with added Gaussian noise, the KF will produce the lowest variance estimate from the underlying data [1], [43]. Internally, a KF achieves this by linearly combining the predictions of an internal model $m$ with the observed data in order to fit a minimum variance normal distribution. Let $Z_{t}=Y_{t}+N_{t}$ be a sequence of data and $M_{t}\sim m(Z_{t-1})$ be a corresponding sequence of model predictions with well defined means and variances. Here, $Y_{t}$ are the true values, $N_{t}$ are i.i.d. Gaussian noise, and $t$ indexing time-points, and $m$ is the internal model of the KF. Then the KF prediction is given by:
$\operatorname{\mathbb{E}}(F_{t})=w_{t}\operatorname{\mathbb{E}}(Z_{t-1})+(1-w_{t})\operatorname{\mathbb{E}}(M_{t})$ (1)
$\operatorname{\sigma^{2}}(F_{t})=w_{t}^{2}\operatorname{\sigma^{2}}(Z_{t-1})+(1-w_{t})^{2}(\operatorname{\sigma^{2}}(M_{t})+Q)$ (2)
$w_{t}=\underset{w^{\prime}}{\textrm{argmin}}[w_{t}^{\prime 2}\operatorname{\sigma^{2}}(Z_{t-1})$
$+(1-w_{t}^{\prime})^{2}(\operatorname{\sigma^{2}}(M_{t})+Q)]$ (3)
where, the terms are as described in 1. $\operatorname{\mathbb{E}}(F_{t})$ denotes the filter estimate of $Y_{t}$, which, in standard KFs is the minimum variance estimate of the process $Y_{t}$. The process uncertainty, $Q$, is traditionally used as a parameter to bias the relative weight between model and observations. Finally the weight $w_{t}$ is the Kalman gain which is monotonically decreasing in $t$ provided $Q$ and $\operatorname{\sigma^{2}}(Z_{t})$ are constant [1]. In the case of a linear model $m$, this monotonicity gives rise to the convergence of the filter variance:
$\operatorname{\sigma^{2}}(F_{t})=\operatorname{\sigma^{2}}(F_{t-1})(1-w_{t})\quad\textrm{with}\quad w_{t}\leq w_{t-1}\leq 1$ (4)
The Kalman gain represents new information gained with each subsequent measurement [43]. A standard KF is depicted as a graphical model in Figure (2A).
GENERAL
$\operatorname{\mathbb{E}}$ Expectation value
$\operatorname{\sigma^{2}}$ Variance
STANDARD KALMAN FILTER (KF)
(subscript $t$ indicates time)
$M_{t}$ Model prediction
$Z_{t}$ Measurement values
$F_{t}$ Filter output
$R$ Data Variance (usually a constant)
$Q$ Process uncertainty (usually a constant)
$w_{t}$ Kalman Gain
$m$ Internal model of the KF
PATHSPACE KALMAN FILTER(PKF)
(subscript $t$ indicates time, superscript $i$ indicates iteration)
$\operatorname{\boldsymbol{M}}^{i}_{t}$ Model prediction for PKF
$\operatorname{\boldsymbol{F}}^{i}_{t}$ Pathspace Filter output
$\operatorname{\sigma^{2}}(Z_{t})$ Dynamic data variance
$\operatorname{\boldsymbol{Q}}^{i}_{t}$ Dynamic Process uncertainty
$\operatorname{\boldsymbol{w}}^{i}_{t}$ Data Weight
$\operatorname{\boldsymbol{v}}^{i}_{t}$ Model Weight
$\operatorname{\boldsymbol{u}}^{i}_{t}$ Weight of previous filter iteration
$\operatorname{\boldsymbol{m}}$ Internal model of the PKF
Table 1: Frequently Used Notations
3 Preliminaries
3.1 Pathspace Kalman Filter (PKF)
To motivate the PKF, we first note that the KF only looks forward in time. In many applications, the entire time-series dataset is already available, and it is desirable that all the dynamics are analyzed together. To do this, we define the PKF which can be seen as an adaptation of the KF where the entire filter iterates over its dynamical output (the path) repeatedly. Additionally, the PKF dynamically updates process uncertainty in order to re-weight the model confidence. Specifically the update equations are:
$\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{F}}_{t}^{i})=\operatorname{\boldsymbol{w}}^{i}_{t}\operatorname{\mathbb{E}}(Z_{t})+\operatorname{\boldsymbol{v}}_{t}^{i}\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{M}}^{i}_{t})+\operatorname{\boldsymbol{u}}_{t}^{i}\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{F}}^{i-1}_{t}),$ (5)
$\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}_{t}^{i})=(\operatorname{\boldsymbol{w}}_{t}^{i})^{2}\operatorname{\sigma^{2}}(Z_{t})$
$+(\operatorname{\boldsymbol{v}}_{t}^{i})^{2}(\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}^{i}_{t})+\operatorname{\boldsymbol{Q}}_{t}^{i-1})+(\operatorname{\boldsymbol{u}}_{t}^{i})^{2}\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}^{i-1}_{t}),$ (6)
$\operatorname{\boldsymbol{w}}^{i}_{t},\operatorname{\boldsymbol{v}}^{i}_{t}=\underset{\operatorname{\boldsymbol{w}}^{\prime i}_{t},\operatorname{\boldsymbol{v}}^{\prime i}_{t}}{\textrm{argmin}}[(\operatorname{\boldsymbol{w}}^{\prime i}_{t})^{2}\operatorname{\sigma^{2}}(Z_{t})$
$+(\operatorname{\boldsymbol{v}}^{\prime i}_{t})^{2}(\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}^{i}_{t})+\operatorname{\boldsymbol{Q}}_{t}^{i-1})+(\operatorname{\boldsymbol{u}}^{\prime i}_{t})^{2}\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}^{i-1}_{t})],$
$\operatorname{\boldsymbol{u}}^{i}_{t}=1-\operatorname{\boldsymbol{w}}_{t}^{i}-\operatorname{\boldsymbol{v}}_{t}^{i},$ (7)
$\operatorname{\boldsymbol{Q}}_{t}^{i}=\operatorname{\boldsymbol{Q}}_{t}^{i-1}+(\operatorname{\boldsymbol{w}}_{t}^{i}+\operatorname{\boldsymbol{v}}_{t}^{i})(L(\operatorname{\boldsymbol{M}}^{i}_{t},Z_{t})-\operatorname{\boldsymbol{Q}}_{t}^{i-1}).$ (8)
There are three different weights $\operatorname{\boldsymbol{w}}$, $\operatorname{\boldsymbol{v}}$ and $\operatorname{\boldsymbol{u}}$ for the data, model, and filter, respectively. The symbols have been made bold to denote that they are vectors in both time $t$ (subscript) and filter iteration $i$ (superscript). For example $\operatorname{\boldsymbol{F}}^{i}=(\operatorname{\boldsymbol{F}}^{i}_{0},...,\operatorname{\boldsymbol{F}}^{i}_{T})$ with $T$ the number of time-points. Notice that the model samples $\operatorname{\boldsymbol{M}}^{i}_{t}\sim\operatorname{\boldsymbol{m}}(\operatorname{\boldsymbol{F}}^{i-1},t)$ are now a function of the same time $t$ and the entire trajectory $\operatorname{\boldsymbol{F}}^{i-1}$ from the previous filter iteration instead of the single filter value $F_{t-1}$ at the previous time. Similarly, the process uncertainty $\operatorname{\boldsymbol{Q}}_{t}^{i-1}$ is now a function of time and iteration instead of a constant. instead of $Q_{t-1}$. Feeding the filter output back into itself is important to ensure the convergence of the PKF as we prove in Appendix 8.1.2. The PKF is illustrated as a graphical model in Figure (2B).
This entire construction can be viewed as an instantiation of the expectation maximization algorithm where in expectation step $i-1$ the filter output $\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{F}}^{i-1})$ is computed. This value is then fed into the maximization step $i$ via the model $\operatorname{\boldsymbol{M}}^{i}\sim\operatorname{\boldsymbol{m}}(\operatorname{\boldsymbol{F}}^{i-1})$ where the filter variance is minimized to compute $\operatorname{\boldsymbol{w}}_{t}^{i}$. The new Kalman gain can then be used to compute the filter output $\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{F}}_{t}^{i})$ and so on.
In the KF formalism, the Kalman gain $\operatorname{\boldsymbol{w}}_{t}$ can be interpreted as the amount of new information gained from the most recent measurement. This value decreases monotonically for a standard KF, meaning that the weighting of the uncertainty associated with the data can never increase in time [43]. Monotonicity of the Kalman gain in time is undesirable because it does not allow for dynamically varying (both increasing and decreasing) the relative contributions to uncertainties from the data and model over time. Next, we prove that for the PKF the Kalman Gain $\operatorname{\boldsymbol{w}}^{i}_{t}$ is monotonically decreasing in $i$ but not $t$.
Theorem 1: The temporal sequence of Kalman Gains in the PKF, $\{\operatorname{\boldsymbol{w}}^{i}_{1},\operatorname{\boldsymbol{w}}^{i}_{2},...,\operatorname{\boldsymbol{w}}^{i}_{t}\}$ for any filter iteration $i$, as defined in equation (7) and updated according equations (5 - 8) may be non-monotonic in time $t$.
Proof Outline: [Detailed proof in Appendix 8.1.1]. For a given filter iteration, $i$, the Kalman Gain ($\operatorname{\boldsymbol{w}}^{i}_{t}$) is given by:
$\operatorname{\boldsymbol{w}}^{i}_{t}=\frac{AB}{AB+BC+CA}$ (9)
where, $A=\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}^{i-1}_{t})$, $B=(\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}^{i}_{t})+\operatorname{\boldsymbol{Q}}^{i-1})$, and $C=\operatorname{\sigma^{2}}(Z)$. As $\operatorname{\boldsymbol{Q}}_{t}^{i-1}$ and $\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}_{t}^{i})$ are not restricted to be monotonic in time, it follows that the sequence of weights, $\operatorname{\boldsymbol{w}}^{i}_{1}$, $\operatorname{\boldsymbol{w}}^{i}_{2}$,…, $\operatorname{\boldsymbol{w}}^{i}_{t}$} may be non-monotonic in time.
Lemma of Theorem 1: The sequence of process uncertainties $\{\operatorname{\boldsymbol{Q}}_{1}^{i},\operatorname{\boldsymbol{Q}}_{2}^{i},...,\operatorname{\boldsymbol{Q}}_{t}^{i}\}$ in time may be non-monotonic in time $t$.
Proof Outline: [Detailed proof in Appendix 8.1.2]. The non-monotonicity of $\operatorname{\boldsymbol{Q}}^{i}_{t}$ originates in the non-monotonicity of the loss function $L$ referenced in Equation (8) when applied to dynamical data. The fact that the process uncertainty can vary over the course of a trajectory allows $\operatorname{\boldsymbol{Q}}^{i}_{t}$ to be used as a dynamic metric to assess the relative sources of uncertainty from the model and data.
Next, we prove that the PKF algorithm converges to a fixed point in iteration space.
Theorem 2: In the limit of infinite iterations, the PKF converges to a fixed point of the process uncertainty $\operatorname{\boldsymbol{Q}}^{*}=\lim_{i\to\infty}\operatorname{\boldsymbol{Q}}^{i}$.
Proof Outline: [Detailed proof in Appendix 8.1.2]. First, we show that the filter variance is monotonically decreasing towards its fixed point $\lim_{i\to\infty}\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}_{t}^{i})$. Then, we show the existence of a fixed point for $\lim_{i\to\infty}\operatorname{\boldsymbol{Q}}_{t}^{i}$ and show that convergence of the filter variance guarantees that the process uncertainty also converges.
Figure 2: Graphical model illustrations of Kalman Filter architectures. Solid arrows indicate dependencies of means and variances. Dashed arrows indicate dependencies only of the variance. A. Graphical Model of a standard Kalman Filter (KF). B. Graphical Model of the path-space Kalman Filter (PKF). Notice that the PKF feeds into itself using the index $i$ instead of in $t$ allowing for an entire trajectory to be iteratively processed.
4 Efficient Bayesian Model Computations
For the KF, it is necessary that there be an internal model $m$ which generates a prediction at the next time-point $t$, $M_{t}\sim m(Z_{t-1})$ with defined mean and $\operatorname{\mathbb{E}}(M_{t})$ and variance $\operatorname{\sigma^{2}}(M_{t})$. Similarly, for the PKF to function, there must be an internal model $\operatorname{\boldsymbol{m}}$ which generates a predictions of the next filter iteration $i$ at each time-point $t$: $\operatorname{\boldsymbol{M}}_{t}^{i}=\operatorname{\boldsymbol{m}}(\operatorname{\boldsymbol{F}}^{i-1},t)$. Notice that in the PKF the entire trajectory can be used to generate the model prediction. In order to expressly represent dynamics, we frame this as a Bayesian inference problem over the parameters of an ordinary differential equation. Consider the following ODE:
$\frac{\textrm{d}x}{\textrm{d}t}=g(x,k)t$ (10)
with solution over a time window $\Delta t=(t_{0},t_{f})$:
$x(k,\Delta t)=\int_{t_{0}}^{t_{f}}g(x;k)\textrm{d}t$ (11)
We define $\operatorname{\boldsymbol{m}}(\zeta,t)=s(X,\Delta t,k)$ as a distribution that assigns solutions of the ODE (10) with parameter values $k$ a probability based on how well the solution fits all the data in a time window $\Delta t=(t_{0},t_{f})$ with $t_{0}\leq t\leq t_{f}$. Following common practice we let:
$s(\zeta,\Delta t,k)=\mathbb{P}(k\mid\zeta)\propto e^{-\mathcal{L}(x(k,\Delta_{t}),X_{\Delta t})}\pi(k)\$ (12)
Here, $\pi(k)$ is a prior over the parameters, $\zeta_{\Delta t}$ denotes a vector of inputs in the range $(t_{0},t_{f})$ (these may come from the Filter $\operatorname{\boldsymbol{F}}$ or the data $Z$), and $\mathcal{L}$ is a loss function (such as the $L_{2}$ norm). In general, solving (12) for high dimensions of $k$ can be numerically challenging. We simplify this problem by choosing $\Delta t$ to be a window of 3 consecutive data points and using carefully chosen two parameter models $g$. This allows us to derive an analytic function $k_{2}=h(k_{1})$ constraining the parameters $k_{1}$ and $k_{2}$ of the model $g$ so it fits the first and third time-points exactly. We then varying a single parameter $k_{1}$ (which unique determines $k_{2}$ via $h$) to generate the distribution $s$ at the middle time-point $t$. This procedure is illustrated graphically in Figure (3). Intuitively $s$ can be thought of as a local distribution of splines consistent with the ODE (11). We describe this procedure more carefully in Appendix 8.3, including a modification necessary for start and end points of the trajectory. Different choices of $g$ are used for different kinds of data as described in the experiments section.
Figure 3: Models are fit across time windows of three measurements, where the first ($t_{0})$ and third data-points ($t_{1})$ are fixed, and the middle point (denoted by the yellow X mark at $t$) is allowed to vary. A. Illustrates the model fit and the process uncertainties when the model variance is relatively large, as represented by the spread of the colored lines, and the process uncertainty is small, as represented by the proximity of the estimated data-point (yellow X mark) to the measured data-point (black dot). B. Illustrates the fit where the process uncertainty is high, but the model variance is low. The implications of these relative measures are further elaborated in Table 2.
4.1 Interpretation of the Bayesian Method and Comparison to Process Uncertainty
In practice, the distribution of splines makes it easy to compute $\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{M}}^{i}_{t})$ and $\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}^{i}_{t})$ by sampling $s(\zeta,\Delta t,k)$ via MCMC in high dimensional cases or by scanning over an array of values $k$ to compute the relevant statistics directly in low dimensional cases. Here, the variance represents how tightly the ODE model can be fit to the input data which is typically the previous filter output $\operatorname{\boldsymbol{F}}^{i-1}$. By contrast, the process uncertainty $\operatorname{\boldsymbol{Q}}_{t}^{i}$ is derived from the error between the model fit $\operatorname{\boldsymbol{M}}_{t}^{i}$ and the underlying data $Z_{t}$. These two sources of uncertainty are illustrated in Figure (3). The total variance at each point of time will be a combination of the variance of the underlying model ($\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{M}}^{i}_{t})$), the process uncertainty ($\operatorname{\boldsymbol{Q}}_{t}^{i}$), and the variance in the data to which the model is fit ($\operatorname{\sigma^{2}}(\operatorname{\boldsymbol{F}}_{t}^{i-1})$). Comparing the process uncertainty, the data variance, and the underlying model allows the PKF to detangle these different sources of uncertainties, which is infeasible using a traditional KF. We illustrate this analysis technique in the following section.
5 Experiments
In order to validate the PKF algorithm, we compared to a real KF using synthetic data because we needed access to the ground-truth state which is typically not present in real datasets. Then, to demonstrate the applicability of the PKF to real world problems, we used it to analyze a gene expression time-course dataset.
5.1 Kalman Filters on Synthetic Data
Data generation: Synthetic population dynamics data was used to compare standard KFs and the PKF models in order to have a ground truth answer to better evaluate the absolute performance of these models. Specifically, data were generated by simulating an ODE with time varying parameters and adding time-varying Gaussian noise to the output:
$\frac{\textrm{d}N}{\textrm{d}t}$ $=N(k_{birth}(t)-k_{death}(t))$ (13)
$Z(t)$ $\sim\mathcal{N}(\mu=N(t),\sigma=k_{noise}(t))$ (14)
with regulation changes represented as time-varying rates:
$k_{birth}(t)$ =$=\begin{cases}0.05&t<5\\ 0.15&t\geq 5\end{cases}\quad k_{death}(t)=\begin{cases}0.05&t<15\\ 0.5&t\geq 15\end{cases}$
and noise injection:
$k_{noise}(t)$ =$=\begin{cases}1.0&t<10\\ 5.0&t\geq 50\end{cases}$
Here $N(t)$ is the ground truth population with time varying birth and death rates, $k_{birth(t)}$ and $k_{death}(t)$. The data $Z(t)$ are $100$ samples per time-point drawn from a normal distribution $\mathcal{N}$ with mean $N(t)$ and time varying standard deviation $k_{noise}(t)$. $Z(t)$ are fed into the the different Kalman Filter models. Mean-square-error is computed by comparing the filter outputs to the ground truth data $N(t)$.
The data generation was done with the objective to demonstrate the ability of the PKF to distinguish between the cases when the internal model is inaccurate with respect to the observed data (this occurs at times when the regulation changes), and when the data is noisy, but the internal model is reliable (which occurs when increasing noise is injected). Specifically, the PKF is able to distinguish the four cases described in Table 2.
Figure 4: Comparison of the KF and PKF on synthetic population dynamics data. A. The simulated dataset consisting of population dynamics generated from simulated a birth-death model and adding Gaussian noise. B. Parameters of the underlying model generating the simulated data. The growth and death rates change at the black vertical lines and the noise increases at the gray vertical line through the end of the time course, which are shown in all figures. C. Filter output a standard KF with $Q=1$. Notice that the filter estimate lags changes in the data generating process. D. Model and data weights for the same KF. E. Process uncertainty, filter variance, and mean-squared error for the same KF. F-G. Filter output and parameters for a standard KF with $Q=10$. Notice that increasing $Q$ decreases model weight. H. Process uncertainty, filter variance, and mean-squared error for the same KF. I. Filter output for the PKF shows no lag. J. Weights for the PKF show that after many iterations, the filter converges to a low variance estimate because the filter weight $\operatorname{\boldsymbol{u}}>\operatorname{\boldsymbol{w}},\operatorname{\boldsymbol{v}}$. K. Mean-square-error, filter variance, and process uncertainty for the PFK. Notice that the mean squared error is lowest for this model. Additionally, the process uncertainty spikes precisely when the data generating process changes.
Results for Synthetic Data: We compared the results of running a standard KF and the PKF on synthetic dataset described above. The internal model used by all the Kalman filter variants also takes the form of a birth-death process, namely the ODE:
$\frac{\textrm{d}x}{\textrm{d}t}=(k_{birth}-k_{death})x.$ (15)
The parameters $k_{birth}$ and $k_{death}$ are fit dynamically over windows of 3 time-points from the underlying data as described in Appendix 8.4.1. To compare the different KF outputs, we use mean-squared-error as a metric:
$\textrm{MSE}=\frac{1}{T}\sum_{t=1}^{T}(\operatorname{\mathbb{E}}(\operatorname{\boldsymbol{F}}_{t}^{i})-N(t))^{2}$ (16)
where $T$ are the total number of timepoints and either a KF or PKF can be used as the filter input into the expected value.
Columns 2 and 3 of Figure (4) illustrate how a standard KF with different process uncertainties fit this dataset. Although increasing the process uncertainty decreases the model weight globally, the traditional KF both still increase the model weight when data becomes noisier. We remark that the Kalman gain (data weight) is not monotonically decreasing because the data variance is a dynamic parameter in this example. Similarly, the filter variance is also non-monotonic in this example due to the use of a non-linear internal model.
The results of the PKF are illustrated in Column 4 of Figure (4). First, we note that the PKF has lower mean-squared-error than any of the other KFs at roughly 0.2% of the KF MSE. Second, we illustrate how the PKF is able to highlight temporal windows when the internal model’s dynamics are insufficient to explain the data. Notice that changes in the data generation parameters as indicated by the black dashed lines result in instantaneous changes in the process uncertainties demonstrating how the PKF can be used to highlight time windows when the internal model cannot fit the data. However, the PKF is able to quickly adapt to these changes and recalibrate its internal model at future time-points. Indeed, even after the noise is increased by a factor of 10 (dashed gray line), the filter is still able to fit the data well. Together, these observations lead to the following heuristic of comparing the process uncertainty $\operatorname{\boldsymbol{Q}}_{t}$ and data variance $\operatorname{\sigma^{2}}({Z_{t}})$ as illustrated in Table (2).
Case (A) in Table 2 represents the ideal situation where the internal model is highly predictive of reliable data. Case (B) in Table 2 represents a failure in the internal model and is an indication that additional temporal resolution or modeling efforts are needed in that time-window. Case (C) in Table 2 indicates that although the underlying data is noisy, the internal model is able to act as a reliable surrogate. Finally, in case (D) in Table 2 it is impossible to determine if the model inaccuracies are due to fitting noise or a lack of understanding about the underlying process.
Low $\operatorname{\boldsymbol{Q}}_{t}$
High $\operatorname{\boldsymbol{Q}}_{t}$
Low $\operatorname{\sigma^{2}}(Z_{t})$
A) Accurate Model
& Reliable Data
B) Inaccurate Model
& Reliable Data
High $\operatorname{\sigma^{2}}(Z_{t})$
C) Accurate Model
& Noisy Data
D) Inaccurate Model
& Noisy Data
Table 2: Interpreting different regimes of process uncertainty $Q_{t}$ and data variance $\operatorname{\sigma^{2}}(Z_{t})$.
Convergence: In Figure (5) we show the numerical convergence of the algorithm on two different time-points. In Appendix (8.1.2) we prove that the PKF converges. This is seen by the rapid convergence of $\operatorname{\boldsymbol{Q}}_{t}^{i}$. The filter variance also converges towards 0, albeit very slowly. This behaviour may seem unexpected, however it occurs due to the feedback of the filter estimates into itself. In the limit of infinite iterations, the PKF has seen its own estimates increasingly more than it has seen the data or the model estimates which gives rise to decreasing filter variance. We emphasize that this behavior is not problematic as the most meaningful quantity in this framework is the process uncertainty and this converges to values representing the dynamic model uncertainty.
Figure 5: Convergence Plots for the PKF. For a time-point in the synthetic data where the Data Variance and Process Uncertainty are both low, Panel A shows Filter Variance and Process Uncertainty and Panel B shows the data, model, and filter weights as a function of iteration. C. For a time-point in the synthetic data where the Data Variance and Process Uncertainty are both high, Panel C shows Filter Variance and Process Uncertainty and Panel D shows the data, model, and filter weights as a function of iteration.
5.2 Application of the Pathspace Kalman Filter to Time-course Gene Expression Data
A prime motivation of this work is its application to biological time-courses, specifically to data obtained through high throughput acquisition methods such as bulk RNAseq and proteomics [23, 12]. In these methods, high dimensional molecular snapshots of biological processes are collected at individual times via destructive experiments. For example, in RNAseq experiments the expression of tens of thousands of genes may be measured simultaneously at each time-point but typically relatively few time-points are collected (usually at a resolution of several hours to days). The high dimensionality of this kind of data coupled with its temporal sparsity raises specific challenges. Firstly, biological systems are dynamic (not static) so one should expect them to change over time [10]. A crucial question is whether the observed changes themselves are expected or unexpected; dynamic process uncertainty allows for the detection of discrepancies between observations and an underlying prior model. Secondly, the scalability of any analysis method to function on high dimensional data is crucial because a central goal of analyzing this kind of data is ascribing novel function to poorly understood genes, for example as targets for therapeutic interventions [37]. In this section, we describe a model of gene expression suitable for detecting biologically important dynamics and apply it to a time-course gene expression dataset.
Dataset Description: In order to demonstrate the scalability of our approach on real world datasets, we used the PKF to analyze time-course gene expression measurements (RNAseq) of a circadian clock dataset in human cancer cell lines [44] with contains of roughly 1.8 million gene expression measurements consisting of 32337 gene expression measurements across two conditions and 14 time-points with two replicates per time-point and condition. Briefly, the mammalian circadian clock is an cell-autonomous oscillatory genetic circuit which regulates diverse cellular functions by synchronizing them to daylight patterns [31], [15]. For a brief description of the mammalian clock, refer to Appendix 8.4.2. In this dataset, the circadian clock is disrupted by activating the gene MyC [44] which causes the clock to be repressed. The effects of repressing the clock can be seen Figure (6A) which shows the dynamics of the core circadian oscillator gene BMAL1. The failure of BMAL1 to oscillate indicates the repression of the circadian clock. Figure (6B) additionally shows the filter output for BMAL1.
Results for Real Gene Expression Data: In this exmaple, an internal model representing gene expression is used:
$\frac{\textrm{d}x}{\textrm{d}t}=k_{exp}-k_{deg}x.$ (17)
Here, the expression rate ($k_{exp}$) and degradation rate ($k_{deg}$) of the transcript are assumed to be constant over short time windows–a hypothesis we call “constant regulation” which we discuss in more detail in Appendix 8.4.2. When the constant regulation model fails to accurately fit the data in a given window, it is indicative that the regulation of the gene is changing and is quantified by higher process uncertainty in the time window. This can be seen for BMAL1 in Figure (6C) where the process uncertainty is generally lower when the circadian clock is repressed.
Figure 6: Application of the PKF to gene expression data in two conditions: Clock Active and Clock Repressed. A. Mean expression data of core circadian clock gene BMAL1 which oscillates when the clock is active and fails to oscillate when it is repressed. The shaded regions denote standard deviation. B. Filter output derived from the data in A. Shaded region denotes filter variance. C. Dynamic Process uncertainty associated with the filter output in B. D. Average log process uncertainty divided by data variance for $1.8$ million gene expression measurements separated by condition and plotted against their percentile variance. Turning off the circadian clock decreases process uncertainty for highly varying genes which is consistent with the circadian clock being a global regulator of gene expression.
Additionally, we ran the PKF on all genes in this dataset and compared the two conditions. The average log ratio of process uncertainty to data variance for these genes is plotted against variance percentile in Figure (6 D). The circadian clock is known to regulate about 20% of the genome [31]. Our analysis picks up this signal by showing increased process uncertainty divided by data variance for the top $50^{\textrm{th}}$ percentile of most varying genes when the circadian clock is active compared to when it is repressed indicating that regulation is less dynamic when the clock is deactivated.
Complexity: Finally, we note that the PKF with efficient Bayesian model computations is a highly scalable method and runs in time linear in the number of time-points, filter iterations, data dimensionality (e.g. number of genes), replicates and experimental conditions. We were able to run it on 1.8 million gene expression measurements on a single 8-core M1-PRO CPU in under an hour by parallelizing the algorithm on a per-gene basis.
6 Related Work
Given their long history, numerous extensions and applications of the KF exist. Famously, KFs have been used in economics [4], parameter estimation, such as estimating insulin sensitivity [2], estimation of electro-chemical charge parameters [19], etc. Here we discuss some general extensions of the KF and their specific use in biology-related applications.
Extensions of the Kalman Filter: Simple extensions include multivariate KFs [5], non-linear or extended KFs [7], unscented KFs [8], and ensembles of KFs [13]. Of particular relevance to our work are Adaptive KFs as derived by Brown and Rutan, which dynamically adapts the filter’s measurement model to the observed data by matching the covariances in the internal model predictions with the covariances of the measurements [3]. Another inspiration for our work, is by Husain et. al., which demonstrates through both theoretical analyses and experiments in photosynthetic cyanobacteria a process of self-tuning the bio-physicial sensitivity to environmental cues, which can be understood within this filtering framework [34].
Recently, there has also been considerable work linking KFs to various machine learning approaches well summarized in a recent review [38]. Several notable examples including using neural networks to fit KF’s process uncertainty and data variance parameters [45], using a neural network as the KF’s internal model [27], and replacing the Gaussian representation of distributions in the KF with normalizing flow neural networks [35], and in hybrid learning models [39].
Methods to Analyze Time-course Biological Data: Numerous methods have been developed to analyze time-course biological datasets. When using gene expression data, differential expression analysis is commonly employed [24]. However, this method does not explicitly take dynamics into account. Several methods have also employed KFs to analyze gene expression time-courses; [28] and [32] use KFs to model gene regulatory networks by detecting statistically significant changes in time-varying linear KFs and using KFs to construct a Bayesian network, respectively. A related approach is to fit a mechanistic model to the dynamical datasets which can be done with non-linear KFs [21] or other methods [40]. However, such approaches typically do not scale to large mechanistic models and suffer from issues related to parameter identifiability [47].
Finally, many methods fit univariate gene expression models RNAseq data [33] and in some cases use dynamically changing fit parameters to model temporal dynamics [41]. However, these models do not directly provide biological insight; instead the task of statistically analyzing the gene expression data is transformed into the related task of statistically analyzing the fit parameters.
7 Conclusion
In summary, our model provides a powerful extension of traditional KFs optimized for time-course datasets with the ability to dynamically highlight temporal regions where the model and data disagree.
Due to the model-driven nature of this work, every new domain requires its own custom internal model development. Although black-box Bayesian methods can be used, these will not scale well and for large data applications custom models must be developed. Currently, we have presented internal models for population dynamics and gene expression. In the future, we hope to add internal models for other biological experiments including single cell sequencing, DNA methylation, CHIPseq, and ATACseq measurements.
In the future, it may be possible to replace custom mechanistic models with black-box AI models. However, one advantage of the domain-specific models is that they require relatively few measurements to run effectively. A fully black-box approach would likely require much larger datasets, which are frequently not available in biological applications.
Finally, we comment that this work only deals with the univariate case which could be extended to a multivariate formalism. Similarly, the assumption of normality of all underlying distributions could potentially be relaxed in future extensions of this method.
Acknowledgements
We thank the El-Samad, Bianco, and Multiscale Modeling groups at Altos Labs for their feedback on this project. We thank Joel Dapello and Carlos Lopez for helpful feedback when preparing this manuscript.